Cheol Soo Lee 9-15 Sinlim Dong Kwan Ak Gu Seoul, Korea

December 16, 2000

Office of Graduate Admission University of Wisonsin-Madison Madison, WI 53706

To whom it may concern;

Enclosed are two copies of my application for admisson to the Department of Statistics. As required, I have sent a copy to the Department. Were any thing missing, please let me know.

Thank you.

Sincerely

Cheol Soo Lee

encl: Application for Admission

cc: The Statistics Department, UWM